# Active Polymer Behavior in Two Dimensions: A Comparative Analysis of Tangential and Push-Pull Models

### v1

@@ I Introduction
Filamentous and polymeric structures are crucial in both biological and artificial systems. They are found across a range of scales, from microscopic components such as microtubules and actin filaments in the cellular cytoskeleton Bray (2000); Schaller et al. (2010a); Ganguly et al. (2012), to bacteria Yaman et al. (2019), worms Deblais et al. (2020a, b); Nguyen et al. (2021), granular systems Wen et al. (2012); Soh et al. (2019), and even macroscopic robotic systems Marvi et al. (2014); Ozkan-Aydin et al. (2021). These structures exhibit a wide range of dynamic behaviors, including self-organization, motility, and responsiveness to external stimuli, which are fundamental to understanding processes like cell division, tissue development, and artificial soft robotics. In recent years, it has become increasingly apparent that the concept of active matter offers a robust framework for understanding the physics of living systems across these diverse scales. Active matter encompasses self-driven systems capable of converting stored or environmental energy into directed motion, leading to collective phenomena. This includes studies of intracellular components Schaller et al. (2010a); Sumino et al. (2012), artificial self-propelled particles Narayan et al. (2007); Theurkauff et al. (2012); Palacci et al. (2013), microorganisms Saw et al. (2017); Be’er and Ariel (2019), large animal species Sumpter (2010); Bialek et al. (2012), robot swarms Rubenstein et al. (2014); Wang et al. (2021), and filamentous matter Isele-Holder et al. (2015); Winkler and Gompper (2020); Philipps et al. (2022); Ubertini et al. (2024); Paolo Malgaretti and Valeriani (2024).
@@
Recently, much attention has been given to the emergent properties of systems combining the activity and conformational degrees of freedom in filamentous and polymeric structures Schaller et al. (2010b); Loi et al. (2011); Suma et al. (2014); Isele-Holder et al. (2015); Eisenstecken et al. (2017); Duman et al. (2018); Prathyusha et al. (2018); Bianco et al. (2018); Shee et al. (2021); Zhu et al. (2024). Known as active polymers Winkler et al. (2017), these systems exhibit dynamics where active forces on individual monomers dictate overall behavior, leading to phenomena such as non-equilibrium phase transitions, dynamic phase separation, and anomalous diffusion Marchetti et al. (2013); Romanczuk et al. (2012); Giomi (2015); Thampi and Yeomans (2016); Alert et al. (2022); Micheletti et al. (2011); Tubiana et al. (2024).
@@
Discrete polar active filaments have been modeled using various representations of active forces Jiang and Hou (2014); Isele-Holder et al. (2015); Bianco et al. (2018), including tangential forces along the polymer backbone Bianco et al. (2018) and push-pull forces Isele-Holder et al. (2015). These models converge to the same continuum limit, making results independent of the discretization method Winkler and Gompper (2020). However, deviations emerge when activity surpasses a threshold, especially in systems with weak bond potentials, large bond-length variations, or small numbers of monomers Jiang and Hou (2014); Isele-Holder et al. (2015). These deviations underscore the need to examine both microscopic details and system dynamics in highly active regimes. Despite these studies, as far as we are aware a direct comparison between tangential and push-pull models is still lacking, and understanding how different active force definitions influence structure and dynamics is crucial, particularly where deviations from continuum models are expected.
@@
In this work, we compare tangential and push-pull models of active filaments in two dimensions, focusing on a version of the push-pull model where the first and last monomers are passive for direct comparison. We analyze these models by varying bending rigidity and polymer length in single-filament systems. As is well known, the system undergoes a transition from open chains to spirals as activity increases Isele-Holder et al. (2015); Prathyusha et al. (2018); Gompper et al. (2020). At low activity, all models exhibit similar structural and dynamic properties. However, as the system approaches the spiral phase, small differences emerge. Nonetheless, in the spiral phase, the behavior converges across all models, regardless of filament length or bending rigidity.
@@
At intermediate densities, we explore phase transitions, including the reentrant phase, previously identified in the push-pull model at high activities  Janzen and Matoz-Fernandez (2024): this phase transitions from spirals back to open chains. We extend this analysis to the tangential and push-pull models with passive head and tail, revealing that while all models exhibit similar phases, the reentrant phase in the tangential model occurs at higher activities due to its smaller active force. Once this force difference is accounted for, the tangential model behaves similarly to the push-pull models at intermediate and high bending rigidities. At low bending rigidity, differences in activity definitions cause slight shifts in the spiral transition. Overall, in two dimensions, once scaled for active force differences, all models behave nearly identically.
@@ II Methods
We simulate a system of active filaments, each consisting of $N_{b}$ beads. We ignore long-range hydrodynamic interactions and consider the dry limit, where damping from the medium dominates, so the surrounding fluid provides only single-particle friction. Thus, the equation of motion for each filament is given by Langevin dynamics,  Suma et al. (2014); Isele-Holder et al. (2015); Prathyusha et al. (2018)
@@
where $m_{i}$ is the mass of bead $i$, $\vec{r}_{i}=(x_{i},y_{i})$ represents the bead’s spatial coordinates and the dot denotes the time derivatives, $\gamma$ is the damping coefficient and $\langle\vec{R}_{i}(t)\cdot\vec{R}_{j}(t^{\prime})\rangle=4\gamma k_{B}T\delta_{ij}\delta(t-t^{\prime})$ is a delta-correlated random force with zero mean and variance, where $k_{B}$ is the Boltzmann constant and $T$ the temperature. Moreover, $\vec{f}_{ij}=-\nabla_{i}\phi(r_{ij})$ is the interaction force between beads $i$ and $j$, where $r_{ij}=\left|\vec{r}_{i}-\vec{r}_{j}\right|$ and $\phi$ is the bonded and short-range nonbonded pair potentials $\phi=\phi_{B}+\phi_{NB}$. Bonded interactions $\phi_{B}=\phi_{bond}+\phi_{bend}$ account for both chain stretching, modeled by the Tether bond potential Noguchi and Gompper (2005) and bending modeled with the harmonic angle potential Prathyusha et al. (2018). Next, the nonbonded interactions, $\phi_{NB}$ account for steric repulsion and are modeled with the Weeks-Chandler-Anderson (WCA) potential Weeks et al. (1971). All details regarding the bonded and nonbonded interaction parameters are described in the Supplementary Material SI . Lastly, $\vec{f}^{a}_{i}$ represents the active self-propulsion force on bead $i$. Here, we employ three models to describe active filaments: the tangential model, where activity is applied along the polymer backbone, mimics biological systems where forces are often tangential to filament movement; the push-pull model is well-suited to capture systems where active forces arise from both polymer extremities, as in some synthetic systems; and lastly the push-pull model with passive head and tail.
@@
The Tangential model: The first model we consider is the tangential (T) model, proposed by Bianco et al. Bianco et al. (2018). In this model, polymer activity is introduced by a tangential self-propulsion force applied along the polymer backbone (see Fig. 1 (a)). The active force on each monomer $i$ is defined as:
@@
with $f_{a}$ the magnitude of the active force directed along the vector ${\vec{r}}_{i+1}-{\vec{r}}_{i-1}$, which is parallel to the tangent of the polymer backbone Bianco et al. (2018); Foglino et al. (2019), and $\vec{t}_{i-1,i+1}$ the unit tangent vector. In this model, the first and last beads of the polymer are passive, meaning that no active force acts on them.
Push-pull model: The second model we examine is the well-known push-pull (PP) model Isele-Holder et al. (2015); Prathyusha et al. (2018) (see Fig. 1(b)), where the active force $\vec{f}^{a}_{i}$ on bead $i$ is given by
@@
Push-pull with passive head and tail: Finally, we introduce the push-pull model with passive head and tail (PHT), which is similar to the standard push-pull model, except the first and last beads are passive. Since the T-model also features passive end beads, this variant allows for a more direct comparison between the tangential and push-pull models
@@
In this study, we focus on polymers with one degree of polymerization $N_{b}=50$, with a monomer size $\sigma=1.0$, interaction energy scale $\epsilon=1.0$, thermal energy $k_{B}T=10^{-1}$, and a damping coefficient $\gamma=1.0$. The length of each active polymer is calculated as $L\approx b\,(N_{b}-1)\sigma$, where $b=0.86$. Note that all parameters are given in LJ units. We define two key dimensionless numbers for our analysis: the scaled persistence length $\xi_{p}/L=2b\kappa/(Lk_{B}T)$, representing the filament’s thermal (passive) stiffness, and the active Péclet number $Pe=f_{a}L^{2}/(\sigma k_{B}T)$, which characterizes the balance between active forces and thermal fluctuations. The bending rigidity $\kappa$ is held constant by setting $\xi_{p}/L=0.06,0.3,1.4$. To investigate filaments in suspension, we set the packing fraction to $\rho=N_{f}N_{b}\,\pi\,\sigma^{2}/4L_{\text{box}}^{2}=0.4$, with the number of filaments $N_{f}=10^{3}$, and varied the Péclet number $Pe$. Additionally, we explored the single-filament case with polymerization degrees $N_{b}=5$ and $N_{b}=15$; the results for these cases, along with further details, are provided in the Supplementary Material SI .
Simulations were performed using the GPU-accelerated SAMOS molecular dynamics package Sknepnek (2024), employing the BAOAB Langevin scheme for numerical integration Leimkuhler and Matthews (2015). Data analysis was carried out with custom Python scripts, and visualizations were generated using ParaView Fabian et al. (2011).
Structural and Dynamical Properties
To characterize the behavior of the three models as a function of the Péclet number and to compare them, we compute the turning number Krantz (1999); Shee et al. (2021), denoted as $\psi$. This quantity is defined as:
@@
where $x$ and $y$ represent the coordinates of the beads, and the primes denote first and second derivatives, respectively. All these properties are averaged over $10^{2}$ distinct equilibrated snapshots and over all the filaments $N_{f}=10^{3}$.
Finally, to effectively compare the different models, we also compute a dynamical property, the mean-squared displacement of the filament’s center of mass, denoted as MSD$(t)$, after ensuring the system is equilibrated,
@@
III Results and discussion
Single filament
We begin comparing the three models by examining the structural properties of the systems in the single-filament case. Figure 2 shows the turning number and curvature for a single filament with $N_{b}=50$ at three different bending rigidities: $\xi_{p}/L=0.06,0.3,$ and $1.4$.
For both structural properties, the three models exhibit the same qualitative behavior. At low $Pe$, the polymers form open chains, and as activity increases, the filaments transition into spirals. Since higher turning number and curvature values correspond to more compact structures, we observe that before the spiral transition, the PHT model shows an slightly higher degree of compactness compared to the other two models, while the T model has the least compact structures. However, once the filaments enter the spiral phase ($\psi>3$), all three models behave similarly. Finally, we observed similar behavior for shorter polymer lengths (see Supplementary Material SI ).
To properly compare the three models, it is important to study not only their structural properties but also their dynamics. Thus, we computed the mean-squared displacement (MSD) for two $Pe$ values, one in the open chain phase ($Pe=10^{3}$) and one in the spiral phase ($Pe=10^{4}$), to assess their dynamic behavior.
Figure 3: Mean squared displacement MSD(t) as a function of time for the three models: blue line represent the T model, red line for the PP model, and green line for the PHT model. (a), (b) and (c) MSD(t) at $Pe=10^{3}$, for $\xi_{p}/L=0.06$, $\xi_{p}/L=0.3$ and $\xi_{p}/L=1.4$, respectively. The mean-squared displacement of the three models are identical, regardless of the bending rigidity. (d), (e) and (f) MSD(t) at $Pe=10^{4}$, for $\xi_{p}/L=0.06$, $\xi_{p}/L=0.3$ and $\xi_{p}/L=1.4$, respectively. In the spiral phase, for small and intermediate bending rigidity ($\xi_{p}/L=0.06,0.3$), the MSD of the T and PP models are nearly identical, while the PHT model is slightly slower. However, all three models share the same slope in their MSD. For large bending rigidity ($\xi_{p}/L=1.4$), the PP and PHT models exhibit the same dynamic behavior, whereas the T model behaves differently, with faster dynamics and a distinct MSD slope compared to the other two models.
@@ III Results and discussion
Figures 4(a) and (d) show the turning number and curvature, respectively, for a bending rigidity of $\xi_{p}/L=0.06$. At this bending rigidity, all three models exhibit very similar behavior. However, the push-pull model with a passive head and tail displays slight deviations from the other two models around $Pe=10^{4}$, due to a higher number of spirals in the system compared to the PP and T models. When the Péclet number increases to approximately $Pe=2.5\times 10^{4}$, all filaments form spirals across all three models. For even higher $Pe$ values ($Pe>2.5\times 10^{4}$), the two push-pull models exhibit the same behavior: the turning number begins to decrease as some spirals start to open, indicating the onset of a reentrant phase. In contrast, the reentrant phase in the tangential model occurs at even higher Péclet numbers. Moreover, as shown in Figure 4(d), curvature changes only slightly with increasing Péclet number. This minimal variation occurs because the small bending allows the filaments to remain highly flexible, even at low Péclet numbers.
@@
Figures 4(b) and (e) represent the turning number and curvature for $\xi_{p}/L=0.3$, respectively. The results show that, for all three models, the turning number increases with increasing Péclet number and then decreases again at very high $Pe$ (i.e., $Pe>10^{4}$). At $Pe\approx 10^{4}$, all filaments form spirals in the system across all models, though the spirals in the THP model are slightly more compact than in the other two models. All three models exhibit a non-monotonic behavior in $\psi$, characteristic of a reentrant phase Janzen and Matoz-Fernandez (2024). However, the reentrant phase occurs at different $Pe$ values for each model. In the T and PHT models, spirals open up at larger $Pe$ compared to the PP model, where the spirals begin to open at $Pe\approx 2.5\times 10^{4}$. Although both the T and PHT models experience a shift in the reentrant phase relative to the PP model, the phase occurs at even higher $Pe$ values in the PHT model than in the T model. As discussed in the previous section (Fig. 3(e)), the dynamics in the spiral phase for the PHT model is slower than in the other two models, leading to less frequent movement and collisions of the spirals. This slower dynamic may explain why a higher $Pe$ is needed in the PHT model to open the spirals and observe a reentrant phase.
@@
Additionally, the curvature (Figure 4(e) ) in the T model is slightly lower than in the other two models, which could be attributed to the different definition of the active force. This variation affects the angle distribution between three consecutive monomers. Overall, the curvature behaves similarly across all three models as a function of the Péclet number: as $Pe$ increases, the curvature also increases, indicating that the filaments become more compact.
@@
Finally, Fig. 4 (c) and (f) represent the turning number and curvature for $\xi_{p}/L=1.4$, respectively. Both panels show that, although all three models exhibit the same qualitative behavior in both turning number and curvature, the pure spiral phase (where all filaments in the system form spirals) and the reentrant phase in the T model are shifted by approximately a factor of two compared to the PP and PHT models.
@@
In summary, these results indicate that the three models share the same qualitative behavior: starting with an open chains phase, followed by a pure spiral phase, and finally a reentrant phase. Consistent with findings at higher density by Prathyusha et al. Prathyusha et al. (2018), the spiral transition occurs at higher $Pe$ for both small and large bending rigidities ($\xi_{p}/L=0.06,1.4$) compared to intermediate bending rigidity ($\xi_{p}/L=0.3$). The only difference between the models is quantitative, as the $Pe$ values at which the spiral and reentrant phases occur are shifted.
@@
Since the active force in the tangential model is defined differently from the push-pull models (PP and PHT), we compute the mean force on each bead to better understand the differences between these models. Given that the spiral transition is driven by a competition between the active and the bending force Prathyusha et al. (2018), Fig. 5 (a), (b), and (c) show the mean of the combined active and bending forces for the T and PHT models, normalized by the mean force in the PP model. This is plotted as a function of $Pe$ for $\xi_{p}/L=0.06,0.3,$ and $1.4$, plotted in the top, middle and bottom panel respectively.

### v2

@@ I Introduction
Filamentous and polymeric structures are crucial in both biological and artificial systems. They are found across a range of scales, from microscopic components such as microtubules and actin filaments in the cellular cytoskeleton bray2000cell ; Schaller2010 ; ganguly2012cytoplasmic , to bacteria Yaman2019 , worms deblais2020phase ; deblais2020rheology ; nguyen2021emergent , granular systems wen2012polymerlike ; soh2019self , and even macroscopic robotic systems marvi2014sidewinding ; ozkan2021collective . These structures exhibit a wide range of dynamic behaviors, including self-organization, motility, and responsiveness to external stimuli, which are fundamental to understanding processes like cell division, tissue development, and artificial soft robotics. In recent years, it has become increasingly apparent that the concept of active matter offers a robust framework for understanding the physics of living systems across these diverse scales. Active matter encompasses self-driven systems capable of converting stored or environmental energy into directed motion, leading to collective phenomena. This includes studies of intracellular components Schaller2010 ; Sumino2012 , artificial self-propelled particles Narayan2007 ; Theurkauff2012 ; Palacci2013 , microorganisms Saw2017 ; Beer2019 , large animal species Sumpter2010 ; Bialek2012 , robot swarms Rubenstein2014 ; Wang2021 , and filamentous matter IseleHolder2015 ; Winkler2020 ; philipps2022tangentially ; Deblais2023 ; ubertini2024universal ; malgaretti2024coil .
@@
Recently, much attention has been given to the emergent properties of systems combining the activity and conformational degrees of freedom in filamentous and polymeric structures schaller2010polar ; Loi2011 ; suma2014motility ; Huijun2014A ; HuijunB ; jiang2014motion ; IseleHolder2015 ; eisenstecken2017internal ; duman2018collective ; Prathyusha2018 ; Bianco2108 ; Shee2021 ; Prathyusha2022 ; zhu_non-equilibrium_2024 . Known as active polymers winkler2017active , these systems exhibit dynamics where active forces on individual monomers dictate overall behavior, leading to phenomena such as non-equilibrium phase transitions, dynamic phase separation, and anomalous diffusion MarchettiRev ; romanczuk2012active ; PhysRevX.5.031003 ; thampi2016active ; alert2022active ; MICHELETTI20111 ; TUBIANA20241 .
@@
Discrete polar active filaments have been modeled using various representations of active forces jiang2014motion ; IseleHolder2015 ; Bianco2108 , including tangential forces along the polymer backbone Bianco2108 and push-pull forces IseleHolder2015 . These models converge to the same continuum limit, making results independent of the discretization method Winkler2020 . However, deviations may emerge when activity surpasses a threshold, especially in systems with weak bond potentials, large bond-length variations, or small numbers of monomers. These deviations underscore the need to examine both microscopic details and system dynamics in highly active regimes. Despite these studies, as far as we are aware a direct comparison between tangential and push-pull models is still lacking, and understanding how different active force definitions influence structure and dynamics is crucial, particularly where deviations from continuum models are expected.
@@
In this work, we compare tangential and push-pull models of active filaments in two dimensions, introducing a version of the push-pull where the first and last monomers are passive to bridge between the two. We analyze these models varying bending rigidity and polymer length in single-filament systems as well as at finite density. As it is well known, at infinite dilution the system undergoes a transition from open chains to spirals as activity increases IseleHolder2015 ; Prathyusha2018 ; gompper20202020 . We find that at low activity, all models exhibit similar structural and dynamic properties. However, as the system approaches the spiral phase, small differences emerge. Nonetheless, in the spiral phase, the behavior converges across all models, regardless of filament length or bending rigidity.
@@
At intermediate densities, we explore phase transitions, including the reentrant phase, previously identified in the push-pull model at high activities janzen2024density : this phase transitions from spirals back to open chains. We extend this analysis to the tangential and push-pull models with passive head and tail, revealing that while all models exhibit similar phases, the reentrant phase in the tangential model occurs at higher activities due to its smaller active force. Once this force difference is accounted for, the tangential model behaves similarly to the push-pull models at intermediate and high bending rigidities. At low bending rigidity, differences in activity definitions cause slight shifts in the spiral transition. Overall, in two dimensions, once scaled for active force differences, all models behave nearly identically.
@@ II Methods
We simulate a system of active filaments, each consisting of $N_{b}$ beads. We ignore long-range hydrodynamic interactions and consider the dry limit, where damping from the medium dominates, so the surrounding fluid provides only single-particle friction. Thus, the equation of motion for each filament is given by Langevin dynamics,  suma2014motility ; IseleHolder2015 ; Prathyusha2018
@@
where $m_{i}$ is the mass of bead $i$, $\vec{r}_{i}=(x_{i},y_{i})$ represents the bead’s spatial coordinates and the dot denotes the time derivatives, $\gamma$ is the damping coefficient and $\langle\vec{R}_{i}(t)\cdot\vec{R}_{j}(t^{\prime})\rangle=4\gamma k_{B}T\delta_{ij}\delta(t-t^{\prime})$ is a delta-correlated random force with zero mean and variance, where $k_{B}$ is the Boltzmann constant and $T$ the temperature. Moreover, $\vec{f}_{ij}=-\nabla_{i}\phi(r_{ij})$ is the interaction force between beads $i$ and $j$, where $r_{ij}=\left|\vec{r}_{i}-\vec{r}_{j}\right|$ and $\phi$ is the bonded and short-range nonbonded pair potentials $\phi=\phi_{B}+\phi_{NB}$. Bonded interactions $\phi_{B}=\phi_{bond}+\phi_{bend}$ account for both chain stretching, modeled by the Tether bond potential Noguchi2005 and bending modeled with the harmonic angle potential Prathyusha2018 . Next, the nonbonded interactions, $\phi_{NB}$ account for steric repulsion and are modeled with the Weeks-Chandler-Anderson (WCA) potential weeks1971role . All details regarding the bonded and nonbonded interaction parameters are described in the Supplementary Material SI . Lastly, $\vec{f}^{a}_{i}$ represents the active self-propulsion force on bead $i$.
In this study, we focus on employing two models of active polymers: the tangential model and the push-pull model. While both models incorporate tangential propulsion, they differ fundamentally in the way active forces are applied along the polymer backbone. The tangential model introduces activity uniformly along the backbone of the polymer, with the active force aligned with the local tangent of the filament. This design reflects biological scenarios where activity is distributed consistently along the filament’s contour. Importantly, in this model, the extremities of the polymer are passive, meaning no active force acts on the first and last monomers. In contrast, the push-pull model applies active forces at each monomer as a combination of contributions from the adjacent bonds. This results in a “push-pull” effect, where activity arises from the interplay of forces on either side of a monomer along the polymer contour. At the extremities, the absence of a neighboring bond naturally reduces the activity at the polymer’s head and tail. This feature makes the push-pull model particularly suited to capture systems where forces originate locally from bond-level interactions.
@@
The Tangential model: The first model we consider is the tangential (T) model, proposed by Bianco et al. Bianco2108 . In this model, polymer activity is introduced by a tangential self-propulsion force applied along the polymer backbone (see Fig. 1 (a)). The active force on each monomer $i$ is defined as:
@@
with $f_{a}$ the magnitude of the active force directed along the vector ${\vec{r}}_{i+1}-{\vec{r}}_{i-1}$, which is parallel to the tangent of the polymer backbone Bianco2108 ; foglino2019 , and $\vec{t}_{i-1,i+1}$ the unit tangent vector. In this model, the first and last beads of the polymer are passive, meaning that no active force acts on them.
Push-pull model: The second model we examine is the well-known push-pull (PP) model IseleHolder2015 ; Prathyusha2018 (see Fig. 1(b)), where the active force $\vec{f}^{a}_{i}$ on bead $i$ is given by
@@
Push-pull with passive head and tail: Finally, we introduce the push-pull model with passive head and tail (PHT), which is similar to the standard push-pull model, except the first and last beads are passive. Since the T-model also features passive end beads, this variant allows to bridge between the tangential and push-pull models.
@@
In this study, we focus on polymers with a degree of polymerization $N_{b}=50$, with a monomer size $\sigma=1.0$, interaction energy scale $\epsilon=1.0$, thermal energy $k_{B}T=10^{-1}$, mass $m=1.0$, and a damping coefficient $\gamma=1.0$, i.e, underdamped limit janzen2024density . The length of each active polymer is calculated as $L\approx b\,(N_{b}-1)\sigma$, where $b=0.86$. Note that all parameters are given in LJ units. We define two key dimensionless numbers for our analysis: the scaled persistence length $\xi_{p}/L=2b\kappa/(Lk_{B}T)$, representing the filament’s thermal (passive) stiffness, and the active Péclet number $Pe=f_{a}L^{2}/(\sigma k_{B}T)$, which characterizes the balance between active forces and thermal fluctuations. The bending rigidity $\kappa$ is held constant by setting $\xi_{p}/L=0.06,0.3,1.4$. To investigate filaments in suspension, we set the packing fraction to $\rho=N_{f}N_{b}\,\pi\,\sigma^{2}/4L_{\text{box}}^{2}=0.4$, with the number of filaments $N_{f}=10^{3}$, and varied the Péclet number $Pe$. Additionally, we explored the single-filament case with polymerization degrees $N_{b}=5$ and $N_{b}=15$; the results for these cases, along with further details, are provided in the Supplementary Material SI .
Simulations were performed using the GPU-accelerated SAMOS molecular dynamics package SAMoS2024 , employing the BAOAB Langevin scheme for numerical integration leimkuhler2015molecular . Data analysis was carried out with custom Python scripts, and visualizations were generated using ParaView fabian2011paraview .
III Results and discussion
Single filament
We begin comparing the three models by examining the structural properties of the systems in the single-filament case. The fist structural property we compute to characterize the behavior of the three models as a function of the Péclet number and to compare them, we compute the turning number krantz1999handbook ; Shee2021 , denoted as $\psi$. This quantity is defined as:
@@ III Results and discussion
where $x$ and $y$ represent the coordinates of the beads, and the primes denote first and second derivatives, respectively. All these properties are averaged over $10^{2}$ distinct decorrelated snapshots at steady state and over all the filaments $N_{f}=10^{3}$.
Figure 2 shows the turning number and curvature for a single filament with $N_{b}=50$ at three different bending rigidities: $\xi_{p}/L=0.06,0.3,$ and $1.4$.
For both structural properties, the three models exhibit the same qualitative behavior. At low $Pe$, the polymers form open chains, and as activity increases, the filaments transition into spirals (see Supplementary Material SI ). Since higher turning number and curvature values correspond to more compact structures, we observe that before the spiral transition, the PHT model shows a slightly higher degree of compactness compared to the other two models, while the T model has the least compact structures. However, once the filaments enter the spiral phase ($\psi>3$), all three models behave similarly. Finally, we observed similar behavior for shorter and longer polymer lengths (see Supplementary Material SI ).
To properly compare the three models, it is important to study not only their structural properties but also their dynamics. Thus, we computed the mean-squared displacement, denoted as MSD$(t)$, after ensuring the system is equilibrated,
@@
Figure 3: Mean squared displacement MSD(t) as a function of time for the three models: blue line represent the T model, red line for the PP model, and green line for the PHT model. (a), (b) and (c) MSD(t) at $Pe=10^{3}$, for $\xi_{p}/L=0.06$, $\xi_{p}/L=0.3$ and $\xi_{p}/L=1.4$, respectively. The mean-squared displacement of the three models are identical, regardless of the bending rigidity.We identify and mark the three regimes observed in the MSD plot as follows: the ballistic regime (denoted as b in the plot) for $t<10^{1}$; superdiffusive (denoted as S in the plot) for $10^{1}<t<10^{3}$ characterized by a slope of approximately 1.5 for $\xi_{p}/L=0.06,0.3$, and 1.8 for $\xi_{p}/L=1.4$; and finally, diffusive (denoted as d in the plot) for $t>10^{3}$. (d), (e) and (f) MSD(t) at $Pe=10^{4}$, for $\xi_{p}/L=0.06$, $\xi_{p}/L=0.3$ and $\xi_{p}/L=1.4$, respectively. For $Pe=10^{4}$, we identify different behaviors depending on the bending. (d) For $\xi_{p}/L=0.06$, similarly to the smaller Péclet ($Pe=10^{3}$), we observe a ballistic regime for $t<10^{0}$, followed by a superdiffusive behavior for $10^{0}<t<10^{2}$ with a slope of approximately 1.5 followed by a diffusive regime for $t>10^{2}$. (e) For $\xi_{p}/L=0.3$, we observe a ballistic regime for $t<10^{0}$, followed by a superdiffusive behavior for $10^{0}<t<10^{2}$ with a slope of approximately 1.4 for the T and PP model. In contrast, we observe a subdiffusive behavior with a slope of 0.8 for the PHT model, followed by a diffusive regime for $t>10^{2}$. (f) For $\xi_{p}/L=1.4$, we observe a ballistic regime for $t<10^{0}$, followed by a superdiffusive behavior for $10^{0}<t<10^{2}$ with a slope of approximately 1.8 for the T model. In contrast, we observe a subdiffusive behavior with a slope of 0.5 for the PP and PHT model, followed by a diffusive regime for $t>10^{2}$.
@@
Figures 4(a) and (d) show the turning number and curvature, respectively, for a bending rigidity of $\xi_{p}/L=0.06$. At this bending rigidity, all three models exhibit very similar behavior. However, the push-pull model with a passive head and tail displays slight deviations from the other two models around $Pe=10^{4}$, due to a higher number of spirals in the system compared to the PP and T models. When the Péclet number increases to approximately $Pe=2.5\times 10^{4}$, all filaments form spirals across all three models. For even higher $Pe$ values ($Pe>2.5\times 10^{4}$), the two push-pull models exhibit the same behavior: the turning number begins to decrease as some spirals start to open, indicating the onset of a reentrant phasejanzen2024density , due to inertia. The reentrant phase occurs also in the T model, but at even higher Péclet numbers. In contrast, as shown in Figure 4(d), the curvature changes only slightly with increasing Péclet number. This minimal variation occurs because the small bending allows the filaments to attain very compact and bent conformations, even at low Péclet numbers.
@@
Figures 4(b) and (e) represent the turning number and curvature for $\xi_{p}/L=0.3$, respectively. The results show that, for all three models, the turning number increases with increasing Péclet number and then decreases again at very high $Pe$ (i.e., $Pe>10^{4}$). At $Pe\approx 10^{4}$, all filaments form spirals in the system across all models, though the spirals in the THP model are slightly more compact than in the other two models. All three models exhibit a non-monotonic behavior in $\psi$, characteristic of the same reentrant phase janzen2024density as before. However, the reentrant phase occurs now at different $Pe$ values for each model. In the T and PHT models, spirals open up at larger $Pe$ compared to the PP model, where the spirals begin to open at $Pe\approx 2.5\times 10^{4}$. Although both the T and PHT models experience a shift in the reentrant phase relative to the PP model, the phase occurs at even higher $Pe$ values in the T model than in the PHT model. As discussed in the previous section (Fig. 3(e)), the dynamics in the spiral phase for isolated chains in the PHT model is slower than in the other two models, leading to less frequent movement and collisions of the spirals. This slower dynamic may explain why a higher $Pe$ is needed, in the PHT model at intermediate values of the bending rigidity, to open the spirals and observe a reentrant phase.
@@
Additionally, the curvature (Figure 4(e) ) in the T model is slightly lower than in the other two models, which could be attributed to the different definition of the active force. This variation affects the angle distribution between three consecutive monomers. Overall, the curvature behaves similarly across all three models as a function of the Péclet number: as $Pe$ increases, the curvature also increases, indicating that the filaments become more compact. The curvature is less sensitive in this context, whereas the turning number clearly emphasizes the reentrant phase.
@@
Finally, Fig. 4 (c) and (f) represent the turning number and curvature for $\xi_{p}/L=1.4$, respectively. Both panels show that, although all three models exhibit the same qualitative behavior in both turning number and curvature, the pure spiral phase (where all filaments in the system form spirals) and the reentrant phase in the T model are shifted by approximately a factor of two compared to the PP and PHT models. It is important to notice, that this factor depends on the bending rigidity as it is clearly shown in Fig. 4(d), and thus, this is not constant.
@@
In summary, these results indicate that the three models share the same qualitative behavior: starting with an open chains phase, followed by a pure spiral phase, and finally a reentrant phase. Consistent with findings at higher density by Prathyusha et al. Prathyusha2018 , in all three models the spiral transition occurs at higher $Pe$ for both small and large bending rigidities ($\xi_{p}/L=0.06,1.4$) compared to intermediate bending rigidity ($\xi_{p}/L=0.3$). The only difference between the models is quantitative, as the $Pe$ values at which the spiral and reentrant phases occur are shifted.
@@
Since the active force in the tangential model is defined differently from the push-pull models (PP and PHT), we compute the mean force over each bead to better understand the differences between these models. Specifically, the spiral transition arises from a competition between the active and bending forces  Prathyusha2018 . Building on this understanding, we focus on the combined contribution of these forces, averaging them across all filaments in the system. Furthermore, to highlight the differences in the forces acting on filaments under the PP model compared to the T and PHT models, we normalize this combined mean force using the values obtained in the PP model. Prathyusha2018 . Figures 5 (a), (b), and (c) show the mean of the combined active and bending forces for the T and PHT models, normalized by the mean force in the PP model. This is plotted as a function of $Pe$ for $\xi_{p}/L=0.06,0.3,$ and $1.4$, plotted in the top, middle and bottom panel respectively.